Find the value of the following:
1.(3t)°
2. 66⁰+10°
3.7x°-(6x)°
4. 100°+100°
5. 4+t°
Thank you

Answer by mananth(16946)


Anything raised to the power 0 is 1
(3t)° =1

2. 66⁰+10°
1+1 =2
3.7x°-(6x)°
7*1-1=6
4. 100°+100°
1+1=2
5. 4+t°
4+1=5